# SGLC: Semantic Graph-Guided Coarse-Fine-Refine Full Loop Closing for LiDAR SLAM

### v1

@@ V Conclusion
This paper presents a semantic graph-guided approach for LiDAR-based loop closing. It initially builds a semantic graph of the foreground instances, which serves as the foundation for descriptor generation and guides subsequent steps in geometric verification and pose estimation. The designed LiDAR scan descriptor leverages both the semantic graph’s topological properties and the appearance characteristics of the background point cloud, thereby enhancing its robustness in loop closure detection. For loop pose estimation, we proposed a coarse-fine-refine registration scheme that considers the alignment of both instance points and background points, offering high efficiency and accuracy. The proposed method supports real-time online operation and is convenient for integration into SLAM systems, particularly semantic SLAM. The experimental results on different datasets demonstrate its superiority.

### v2

@@ V Conclusion
This paper presents a semantic graph-guided approach for LiDAR-based loop closing. It builds a semantic graph of the foreground instances for descriptor generation and guides subsequent geometric verification and pose estimation. The designed LiDAR scan descriptor leverages both the semantic graph’s topological properties and the appearance characteristics of the background points, thereby enhancing its robustness in LCD. For pose estimation, we proposed a coarse-fine-refine registration scheme that considers the alignment of both instance and background points, offering high efficiency and accuracy. The method supports real-time online operation and is convenient for integration into SLAM systems. The experimental results on multiple datasets demonstrate its superiority.